OpenStudy (anonymous):

Would someone help me find my next step in the operation: y^2+2y-8 / y^2 +7y+12 So far, I have simplified the top to (y-2)(y-2). Am I right so far?And what is my next step please? :)

OpenStudy (anonymous):

((y+4)(y-2))/((y+4)(y+3))

OpenStudy (anonymous):

Thank you so much, but is this the answer? Can you tell me how you got there? Sorry to be a bother, I just want to make sure I understand :)

OpenStudy (anonymous):

Oh! This is the next step right? But how did you factor to get there?

OpenStudy (anonymous):

The y+4's would cancel?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you, thank you!! c:

OpenStudy (mathmale):

Strong suggestion: Any time you factor, multiply your factors back together to check whether they produce the original polynomial. Note that (y-2)(y-2) = y^2 - 4x +4, which does not agree with your original denominator.

OpenStudy (anonymous):

So is y-2/y+3 not the correct one? How so?

OpenStudy (anonymous):

y-2/y+3 is currect

OpenStudy (anonymous):

Okay, thank you @scorpions786 :):):)

OpenStudy (anonymous):

@scorpions786 , would you be interested in helping me with another?

OpenStudy (anonymous):

but keep in mind y can't be -4 because we just simplify it with out simplification \[\frac{ x+4 }{ x+4 } . \frac{ x-2 }{ x+3 } \] you see what happen if y=-4

OpenStudy (anonymous):

Would this be a restriction? I've only learned of restrictions that are "x="

OpenStudy (anonymous):

ops it's wrong format exchange x with y

OpenStudy (anonymous):

Okie :) haha, it can't be equal to zero correct? So, this wouldn't work? I'm not quite sure..

OpenStudy (anonymous):

it will be zero if it was -4 ^^

OpenStudy (anonymous):

so yes ^^ correct good luck :D

OpenStudy (anonymous):

Thank you :) Much appreciated! Would you like to help me with another, if you have time to of course :)

OpenStudy (anonymous):

alrighty :)